# Flow matching for reaction pathway generation

### v3

@@ 1 Introduction
There have been several classes of automated methods for generating reaction networks from specified reactants using quantum chemical calculations [1]. Template-based approaches encode known elementary reaction rules to construct networks of potential intermediates and products [2]. Physics-based methods identify transition states to assess kinetic feasibility, followed by intrinsic reaction coordinate analyses to trace reaction pathways [3, 4, 5, 6, 7, 8, 9]. Graph-based algorithms represent molecules as graphs and recursively enumerate intermediates accessible via single-step transformations [10, 11], subsequently linking reactants and products through techniques such as the growing string and nudged elastic band methods. While these strategies provide powerful frameworks for elucidating mechanisms and discovering novel pathways, they remain computationally demanding. However, constructing even a moderately sized reaction network can require thousands of elementary steps, demanding millions of DFT single-point evaluations [12]. This computational bottleneck has spurred the development of machine learning approaches capable of modeling complex chemical transformations with greater efficiency.
Diffusion models have emerged as state-of-the-art approaches for molecular generation, wherein a neural network learns the underlying probability distribution and generates new configurations by modeling a reverse-time stochastic diffusion process. This process gradually transforms a simple prior distribution (e.g., Gaussian noise) into a complex data distribution. Numerous diffusion frameworks have been developed for central problems in physics and chemistry [13, 14, 15]. In mechanistic chemistry, diffusion-based approaches such as TSDiff [16] and OA-ReactDiff [17] have been introduced to generate the transition states of each elementary reaction in a reaction network, replacing the costly nudged elastic band or string methods. However, their reliance on stochastic differential equations (SDEs) introduces limitations familiar to all diffusion approaches: stochastic noise complicates training, slows sampling, and makes conditional generation difficult to control.
Flow matching offers a deterministic alternative. Instead of learning a stochastic reverse diffusion process, a flow-matching model learns the velocity field of an ordinary differential equation (ODE) that continuously transports a prior distribution to a target distribution. This replaces the noisy SDE formulation of diffusion models with a smooth, analytically defined flow, yielding stable training, precise trajectory control, and dramatically faster sampling. Conceptually, flow matching unifies the family of SDE-based diffusion models, rectified flows [18], and consistency models [19] under a single deterministic transport framework.
@@ 3.1 Generating transition states
In the following, we compare the performance of MolGEN with that of diffusion models TSDiff and OA-ReactDiff, as well as a double-ended flow matching model [25] where a custom prior is designed for each target configuration. MolGEN generates a stochastic distribution by a deterministic pass, indicating two key characteristics. (1) The generation occurs near stationary points on the potential energy surface, where the energy gradient approaches zero. This contrasts with the double-ended flow matching model, which yield a unique configuration. Considering this, we generate 30 independent samples for each test TS and select the one with the lowest energy as the final prediction. Since these samples can be generated in parallel, this approach does not necessarily increase inference time. (2) MolGEN is also capable of identifying alternative TS configurations, including those with lower transition barriers than the reference data, an ability that double-ended flow-matching models lack.
@@
    The values are borrowed from Duan [25], who used Transition1x database [20] for training and testing.
@@
    The values are borrowed from Kim [16], who used the database reported by Grambow et al. [26] for training and testing.
MolGEN achieves a mean root mean square deviation (RMSD) of approximately $0.106\ \rm\mathring${A} between the generated and reference TS structures on the Transition1x test set, halving the error of diffusion models and matching the error of the double-ended flow matching model [25] (Table 1). The mean barrier height error produced by MolGEN is about $3.285\rm\ eV$, also halving that of diffusion models and comparable to that of the double-ended flow matching model ($3.608\ \rm eV$). In chemical reactions, a difference of one order of magnitude in reaction rate is often used as the characterization of chemical accuracy, which translates to $1.58\ \rm kcal/mol$ in barrier height error assuming a reaction temperature of 70C70\ $^{\circ}\rm$ C. By this standard, over $54\%$ of TS from MolGEN have high chemical accuracy relative to the reference data, on par with that of the state-of-the-art of the double-ended flow matching model ($57\%$).
@@
The performance of MolGEN can be further enhanced by training on a larger and more diverse dataset. We combined RGD1 [27] with Transition1x for this purpose. RGD1 covers 177,992 reactions involving molecules containing $C$, $H$, NN, and $O$ atoms, with up to 10 heavy atoms. The TSs and products in RGD1 are derived from NEB obtained with DFT (B3LYP-D3/TZVP). Of these, 3% and 6% are held out as a test and validation set, respectively. Training on this expanded dataset reduced the mean RMSD to $0.100\ \rm\mathring${A} and the mean barrier height error to $2.759\ \rm eV$. More notably, the ratio of valid TS increased to $89.20\%$, while the ratio of TS with energies lower than reference rose sharply to $27.87\%$, approaching the performance achieved by the diffusion-based model TSDiff [16] ($30.96\%$).
@@
We can drive the model to seek stationary point even further through training only by KL divergence loss, as evidenced in Table 1. When trained only by KL divergence, the ratio of TS with energies lower than reference rose to $30.66\%$ matching the performance of diffusion models [16], even though the mean RMSD increased slightly to $0.112\ \rm\mathring${A}. This indicates that the model is locating more alternative TS with lower energies than the reference data.
@@
Since the ODE follows an optimal transport trajectory, it theoretically requires only two evaluations to reach the same target state. To quantify the deviation of the learned dynamics from this optimal transport path, we systematically reduce the number of function evaluations (nfe) during sampling. Empirically, MolGEN attains convergence at nfe=10, at which point the mean RMSD of the generated TS stabilizes, corresponding to an inference time of $0.8\ \rm s$. In practice, we employ the dopri5 adaptive integrator [28] to guarantee numerical stability and convergence, yielding an average inference time of $2.1\ \rm s$.
@@ 3.2 Generating reaction products
We assessed predictive performance using top-k accuracy, which quantifies the model’s ability to correctly anticipate plausible products for a given reactant. Leveraging the distributional nature of conditional flow matching, MolGEN can be sampled repeatedly for the same reactant to generate a diverse ensemble of candidate products. Each predicted product geometry is converted to an adjacency matrix [29] and ranked according to its sampling frequency.
@@
We compare MolGEN’s top-k accuracy against that of a prototypical sequence-based model, Graph2SMILES [30], and a sequence model with electron distribution encoding, FlowER [31]. While MolGEN achieves a lower top-1 accuracy (43%) compared to Graph2SMILES (80%), its top-3 (and larger k) accuracy rises sharply to near 100% (Figure 2a), on par with FlowER and surpassing Graph2SMILES, which saturates at 90%. Notably, since MolGEN preserves atomic identities and stoichiometry, it inherently avoids the hallucinatory artifacts often observed in sequence-based models, where mass or electron conservation can be violated.
@@ 3.3 Toward a Self-Contained Solution for Reaction Network Generation
For product generation, traditional reaction network exploration packages such as Yet Another Reaction Program (YARP) [33, 34] employ the graph enumeration to explore potential products based on empirical rules like b2f2 (break 2 bonds and form 2 new bonds). While effective, such rules may overlook feasible products. TS sampling presents an additional challenge. Since reactant-product pairs can be aligned in infinitely many ways, the alignment influences the success rate of subsequent TS refinements. YARP employs heuristic criteria, such as mass-weighted root-mean-square displacement, to identify alignments likely to yield successful TS optimizations via nudged elastic band (NEB) or growing string method (GSM).
@@
The fractions of valid and intended transition-state (TS) structures, along with the average number of gradient descent calls (excluding IRC calculations) required for TS optimization, are summarized in Figure 2d and compared with results from string-based methods reported by Grambow et al. [32]. Both the valid and intended TS fractions achieved by MolGEN exceed those of the string methods. Moreover, the number of gradient descent calls is drastically reduced to 12 per reaction, indicating that the TS geometries generated by MolGEN lie in close proximity to the corresponding TS. The associated energy changes following TS relaxation are shown in Figure 2b.
@@ 5.1.1 Preliminaries: continuous normalizing flows
Chen et al. [35] suggested modeling $\boldsymbol{v}_{t}$ by a neural network $\boldsymbol{v}_{t}^{\theta}(\boldsymbol{x})$, where $\theta$ are learnable parameters, yielding a deep parametrization of the flow $\phi_{t}$, called a continuous Normalizing Flow (CNF). A CNF pushes a simple base density $p_{0}$ (e.g. Gaussian noise) to a complex target $p_{1}$ via
@@ 5.1.2 Flow matching
A simple way to construct a target probability path is via a mixture of simpler conditional paths [36]. Given a data sample $\boldsymbol{x}_{1}$, we denote by $p_{t}(\boldsymbol{x}|\boldsymbol{x}_{1})$ a conditional probability path such that it satisfies $p_{0}(\boldsymbol{x}|\boldsymbol{x}_{1})=p_{0}$ at $t=0$, and we design $p_{1}(\boldsymbol{x}|\boldsymbol{x}_{1})$ to be concentrated around $\boldsymbol{x}_{1}$ (e.g., $p_{1}(\boldsymbol{x}|\boldsymbol{x}_{1})=\mathcal{N}(\boldsymbol{x}|\boldsymbol{x}_{1},\sigma^{2}\boldsymbol{I})$ for small $\sigma>0$). Marginalizing over the unknown data distribution $q(\boldsymbol{x}_{1})$ gives the marginal probability path [36]
@@
One can also define a marginal vector field by “mixing” the conditional fields [36]:
@@ 5.1.3 Gaussian probability path
The flow matching objective works with any choice of probability path and vector field. We can construct $p_{t}(\boldsymbol{x}|\boldsymbol{x}_{1})$ and $\boldsymbol{u}_{t}(\boldsymbol{x}|\boldsymbol{x}_{1})$ for a general family of Gaussian conditional paths [36]. Namely, we set
@@
Among the infinitely many vector fields generating this path, Lipman et al. proposed to choose the simplest affine flow conditioned on $\boldsymbol{x}_{1}$ [36]:
@@
where $\theta$ are learnable parameters of a neural network model. $\mathcal{U}[0,1]$ is a uniform distribution over the interval $[0,1]$. The CFM loss have identical gradients w.r.t. $\theta$ as the FM loss Eq. 1, but is simpler to compute. Thus, it is the common choice for flow matching training [36].
In practice, we use a regression based objective [37]
@@ 5.1.4 Optimal transport path
Then this path is generated by the velocity field [36]
@@ 5.1.6 Dynamic optimal transport
Dynamic optimal transport, as introduced by Benamou and Brenier [38], presents the squared Wasserstein-2 distance $W^{2}_{2}(p_{0},p_{1})$ in a dynamic formulation. This formulation involves optimizing the squared $L^{2}$ norm of a time-varying vector field $\boldsymbol{u}_{t}(\boldsymbol{x})$, which transports samples from $\boldsymbol{x}_{0}\sim p_{0}(\boldsymbol{x}_{0})$ to $\boldsymbol{x}_{1}\sim p_{1}(\boldsymbol{x}_{1})$ according to an ordinary differential equation $\frac{d\psi_{t}(\boldsymbol{x})}{dt}=u_{t}(\boldsymbol{x})$. Mathematically, this dynamic formulation is expressed as
@@
where $\Pi(p_{0},p_{1})$ is the set of couplings with $p_{0}$ and $p_{1}$. This approach is in contrast with the flow matching method by Lipman et al. [36], where the probability path is given by Gaussian distributions. In ReactOT [25], Duan et al. assumed the optimal coupling to be $\pi^{*}\sim(\frac{R+P}{2},TS)$. This enables a flow matching model that generates a unique TS for a given reactant-product pair.
@@ 5.2.2 Construction of edge features
An edge feature $\mathcal{E}=\oplus^{l_{\rm max}}_{l=0}d_{l}l^{p(l)}$ is constructed as $d_{l}$ copies of degree-$l$ spherical components with parity $p(l)=\text{even}$ for even $l$ and odd for odd $l$, so that the edge features transform as proper SO(3) tensors. This mirrors the standard design in higher-order equivariant MPNNs where messages are built by contracting node features with spherical harmonics of edge directions and radial envelopes, then mapped via Clebsch–Gordan (tensor-product) rules to the desired output irreps [39].

### v4

@@ 1 Introduction
There have been several classes of automated methods for generating reaction networks from specified reactants using quantum chemical calculations [dewyer2018methods]. Template-based approaches encode known elementary reaction rules to construct networks of potential intermediates and products [rappoport2014complex]. Physics-based methods identify transition states to assess kinetic feasibility, followed by intrinsic reaction coordinate analyses to trace reaction pathways [ohno2004scaled, maeda2005global, maeda2011finding, maeda2013systematic, garay2022new, bhoorasingh2015transition, martinez2015automated]. Graph-based algorithms represent molecules as graphs and recursively enumerate intermediates accessible via single-step transformations [dewyer2017finding, yang2017automatic], subsequently linking reactants and products through techniques such as the growing string and nudged elastic band methods. While these strategies provide powerful frameworks for elucidating mechanisms and discovering novel pathways, they remain computationally demanding. However, constructing even a moderately sized reaction network can require thousands of elementary steps, demanding millions of DFT single-point evaluations [von2020exploring]. This computational bottleneck has spurred the development of machine learning approaches capable of modeling complex chemical transformations with greater efficiency.
Diffusion models have emerged as state-of-the-art approaches for molecular generation, wherein a neural network learns the underlying probability distribution and generates new configurations by modeling a reverse-time stochastic diffusion process. This process gradually transforms a simple prior distribution (e.g., Gaussian noise) into a complex data distribution. Numerous diffusion frameworks have been developed for central problems in physics and chemistry [xie2021crystal, zeni2023mattergen, jiao2023crystal]. In mechanistic chemistry, diffusion-based approaches such as TSDiff [kim2024diffusion] and OA-ReactDiff [duan2023accurate] have been introduced to generate the transition states of each elementary reaction in a reaction network, replacing the costly nudged elastic band or string methods. However, their reliance on stochastic differential equations (SDEs) introduces limitations familiar to all diffusion approaches: stochastic noise complicates training, slows sampling, and makes conditional generation difficult to control.
Flow matching offers a deterministic alternative. Instead of learning a stochastic reverse diffusion process, a flow-matching model learns the velocity field of an ordinary differential equation (ODE) that continuously transports a prior distribution to a target distribution. This replaces the noisy SDE formulation of diffusion models with a smooth, analytically defined flow, yielding stable training, precise trajectory control, and dramatically faster sampling. Conceptually, flow matching unifies the family of SDE-based diffusion models, rectified flows [liu2022flow], and consistency models [song2023consistency] under a single deterministic transport framework.
@@ 3.1 Generating transition states
In the following, we compare the performance of MolGEN with that of diffusion models TSDiff and OA-ReactDiff, as well as a double-ended flow matching model [duan2025optimal] where a custom prior is designed for each target configuration. MolGEN generates a stochastic distribution by a deterministic pass, indicating two key characteristics. (1) The generation occurs near stationary points on the potential energy surface, where the energy gradient approaches zero. This contrasts with the double-ended flow matching model, which yield a unique configuration. Considering this, we generate 30 independent samples for each test TS and select the one with the lowest energy as the final prediction. Since these samples can be generated in parallel, this approach does not necessarily increase inference time. (2) MolGEN is also capable of identifying alternative TS configurations, including those with lower transition barriers than the reference data, an ability that double-ended flow-matching models lack.
@@
    The values are borrowed from Duan [duan2025optimal], who used Transition1x database [schreiner2022transition1x] for training and testing.
@@
    The values are borrowed from Kim [kim2024diffusion], who used the database reported by Grambow et al. [grambow2020reactants] for training and testing.
MolGEN achieves a mean root mean square deviation (RMSD) of approximately $0.106\ \rm\mathring${A} between the generated and reference TS structures on the Transition1x test set, halving the error of diffusion models and matching the error of the double-ended flow matching model [duan2025optimal] (Table 1). The mean barrier height error produced by MolGEN is about $3.285\rm\ kcal/mol$, also halving that of diffusion models and comparable to that of the double-ended flow matching model ($3.608\ \rm kcal/mol$). In chemical reactions, a difference of one order of magnitude in reaction rate is often used as the characterization of chemical accuracy, which translates to $1.58\ \rm kcal/mol$ in barrier height error assuming a reaction temperature of 70C70\ $^{\circ}\rm$ C. By this standard, over $54\%$ of TS from MolGEN have high chemical accuracy relative to the reference data, on par with that of the state-of-the-art of the double-ended flow matching model ($57\%$).
@@
The performance of MolGEN can be further enhanced by training on a larger and more diverse dataset. We combined RGD1 [zhao2023comprehensive] with Transition1x for this purpose. RGD1 covers 177,992 reactions involving molecules containing $C$, $H$, NN, and $O$ atoms, with up to 10 heavy atoms. The TSs and products in RGD1 are derived from NEB obtained with DFT (B3LYP-D3/TZVP). Of these, 3% and 6% are held out as a test and validation set, respectively. Training on this expanded dataset reduced the mean RMSD to $0.100\ \rm\mathring${A} and the mean barrier height error to $2.759\ \rm kcal/mol$. More notably, the ratio of valid TS increased to $89.20\%$, while the ratio of TS with energies lower than reference rose sharply to $27.87\%$, approaching the performance achieved by the diffusion-based model TSDiff [kim2024diffusion] ($30.96\%$).
@@
We can drive the model to seek stationary point even further through training only by KL divergence loss, as evidenced in Table 1. When trained only by KL divergence, the ratio of TS with energies lower than reference rose to $30.66\%$ matching the performance of diffusion models [kim2024diffusion], even though the mean RMSD increased slightly to $0.112\ \rm\mathring${A}. This indicates that the model is locating more alternative TS with lower energies than the reference data.
@@
Since the ODE follows an optimal transport trajectory, it theoretically requires only two evaluations to reach the same target state. To quantify the deviation of the learned dynamics from this optimal transport path, we systematically reduce the number of function evaluations (nfe) during sampling. Empirically, MolGEN attains convergence at nfe=10, at which point the mean RMSD of the generated TS stabilizes, corresponding to an inference time of $0.8\ \rm s$. In practice, we employ the dopri5 adaptive integrator [hairer1993solving] to guarantee numerical stability and convergence, yielding an average inference time of $2.1\ \rm s$.
@@ 3.2 Generating reaction products
We assessed predictive performance using top-k accuracy, which quantifies the model’s ability to correctly anticipate plausible products for a given reactant. Leveraging the distributional nature of conditional flow matching, MolGEN can be sampled repeatedly for the same reactant to generate a diverse ensemble of candidate products. Each predicted product geometry is converted to an adjacency matrix [seo2021topology] and ranked according to its sampling frequency.
@@
We compare MolGEN’s top-k accuracy against that of a prototypical sequence-based model, Graph2SMILES [tu2022permutation], and a sequence model with electron distribution encoding, FlowER [joung2025electron]. While MolGEN achieves a lower top-1 accuracy (43%) compared to Graph2SMILES (80%), its top-3 (and larger k) accuracy rises sharply to near 100% (Figure 2a), on par with FlowER and surpassing Graph2SMILES, which saturates at 90%. Notably, since MolGEN preserves atomic identities and stoichiometry, it inherently avoids the hallucinatory artifacts often observed in sequence-based models, where mass or electron conservation can be violated.
@@ 3.3 Toward a Self-Contained Solution for Reaction Network Generation
For product generation, traditional reaction network exploration packages such as Yet Another Reaction Program (YARP) [zhao2021simultaneously, zhao2022algorithmic] employ the graph enumeration to explore potential products based on empirical rules like b2f2 (break 2 bonds and form 2 new bonds). While effective, such rules may overlook feasible products. TS sampling presents an additional challenge. Since reactant-product pairs can be aligned in infinitely many ways, the alignment influences the success rate of subsequent TS refinements. YARP employs heuristic criteria, such as mass-weighted root-mean-square displacement, to identify alignments likely to yield successful TS optimizations via nudged elastic band (NEB) or growing string method (GSM).
@@
The fractions of valid and intended transition-state (TS) structures, along with the average number of gradient descent calls (excluding IRC calculations) required for TS optimization, are summarized in Figure 2d and compared with results from string-based methods reported by Grambow et al. [grambow2018unimolecular]. Both the valid and intended TS fractions achieved by MolGEN exceed those of the string methods. Moreover, the number of gradient descent calls is drastically reduced to 12 per reaction, indicating that the TS geometries generated by MolGEN lie in close proximity to the corresponding TS. The associated energy changes following TS relaxation are shown in Figure 2b.
@@ 5.1.1 Preliminaries: continuous normalizing flows
Chen et al. [chen2018neural] suggested modeling $\boldsymbol{v}_{t}$ by a neural network $\boldsymbol{v}_{t}^{\theta}(\boldsymbol{x})$, where $\theta$ are learnable parameters, yielding a deep parametrization of the flow $\phi_{t}$, called a continuous Normalizing Flow (CNF). A CNF pushes a simple base density $p_{0}$ (e.g. Gaussian noise) to a complex target $p_{1}$ via
@@ 5.1.2 Flow matching
A simple way to construct a target probability path is via a mixture of simpler conditional paths [lipman2022flow]. Given a data sample $\boldsymbol{x}_{1}$, we denote by $p_{t}(\boldsymbol{x}|\boldsymbol{x}_{1})$ a conditional probability path such that it satisfies $p_{0}(\boldsymbol{x}|\boldsymbol{x}_{1})=p_{0}$ at $t=0$, and we design $p_{1}(\boldsymbol{x}|\boldsymbol{x}_{1})$ to be concentrated around $\boldsymbol{x}_{1}$ (e.g., $p_{1}(\boldsymbol{x}|\boldsymbol{x}_{1})=\mathcal{N}(\boldsymbol{x}|\boldsymbol{x}_{1},\sigma^{2}\boldsymbol{I})$ for small $\sigma>0$). Marginalizing over the unknown data distribution $q(\boldsymbol{x}_{1})$ gives the marginal probability path [lipman2022flow]
@@
One can also define a marginal vector field by “mixing” the conditional fields [lipman2022flow]:
@@ 5.1.3 Gaussian probability path
The flow matching objective works with any choice of probability path and vector field. We can construct $p_{t}(\boldsymbol{x}|\boldsymbol{x}_{1})$ and $\boldsymbol{u}_{t}(\boldsymbol{x}|\boldsymbol{x}_{1})$ for a general family of Gaussian conditional paths [lipman2022flow]. Namely, we set
@@
Among the infinitely many vector fields generating this path, Lipman et al. proposed to choose the simplest affine flow conditioned on $\boldsymbol{x}_{1}$ [lipman2022flow]:
@@
where $\theta$ are learnable parameters of a neural network model. $\mathcal{U}[0,1]$ is a uniform distribution over the interval $[0,1]$. The CFM loss have identical gradients w.r.t. $\theta$ as the FM loss Eq. 1, but is simpler to compute. Thus, it is the common choice for flow matching training [lipman2022flow].
In practice, we use a regression based objective [albergo2023stochastic]
@@ 5.1.4 Optimal transport path
Then this path is generated by the velocity field [lipman2022flow]
@@ 5.1.6 Dynamic optimal transport
Dynamic optimal transport, as introduced by Benamou and Brenier [benamou2000computational], presents the squared Wasserstein-2 distance $W^{2}_{2}(p_{0},p_{1})$ in a dynamic formulation. This formulation involves optimizing the squared $L^{2}$ norm of a time-varying vector field $\boldsymbol{u}_{t}(\boldsymbol{x})$, which transports samples from $\boldsymbol{x}_{0}\sim p_{0}(\boldsymbol{x}_{0})$ to $\boldsymbol{x}_{1}\sim p_{1}(\boldsymbol{x}_{1})$ according to an ordinary differential equation $\frac{d\psi_{t}(\boldsymbol{x})}{dt}=u_{t}(\boldsymbol{x})$. Mathematically, this dynamic formulation is expressed as
@@
where $\Pi(p_{0},p_{1})$ is the set of couplings with $p_{0}$ and $p_{1}$. This approach is in contrast with the flow matching method by Lipman et al. [lipman2022flow], where the probability path is given by Gaussian distributions. In ReactOT [duan2025optimal], Duan et al. assumed the optimal coupling to be $\pi^{*}\sim(\frac{R+P}{2},TS)$. This enables a flow matching model that generates a unique TS for a given reactant-product pair.
@@ 5.2.2 Construction of edge features
An edge feature $\mathcal{E}=\oplus^{l_{\rm max}}_{l=0}d_{l}l^{p(l)}$ is constructed as $d_{l}$ copies of degree-$l$ spherical components with parity $p(l)=\text{even}$ for even $l$ and odd for odd $l$, so that the edge features transform as proper SO(3) tensors. This mirrors the standard design in higher-order equivariant MPNNs where messages are built by contracting node features with spherical harmonics of edge directions and radial envelopes, then mapped via Clebsch–Gordan (tensor-product) rules to the desired output irreps [batzner20223].